OpenStudy (anonymous):

4x + 1 over 3y2 What is the value of the expression above when x = 2 and y = 3? You must show all work and calculations to receive full credit.

OpenStudy (michele_laino):

substituting your numerical data into your expression, we get: \[4x + \frac{1}{3}{y^2} = 4 \cdot 2 + \frac{1}{3} \cdot {3^2} = ...\]

OpenStudy (michele_laino):

here, we have to compute the power, first

OpenStudy (anonymous):

can u sum it down a little im smart just not that smart

OpenStudy (michele_laino):

ok! next step is: \[4x + \frac{1}{3}{y^2} = 4 \cdot 2 + \frac{1}{3} \cdot {3^2} = \left( {4 \cdot 2} \right) + \left( {\frac{1}{3} \cdot 9} \right) = ...?\]

OpenStudy (anonymous):

um what do i have to do

OpenStudy (michele_laino):

you have to do multiplications first

OpenStudy (anonymous):

ok wht is the factors i use to multiply

OpenStudy (michele_laino):

first step: \(4 \cdot 2=...\)

OpenStudy (anonymous):

8

OpenStudy (michele_laino):

ok! then: \[{\frac{1}{3} \cdot 9}=...?\]

OpenStudy (anonymous):

um 1 x 9 is 9 3 x 9 is 27

OpenStudy (michele_laino):



OpenStudy (anonymous):

so 3 x 1

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

what is \(3 \cdot 1=...?\)

OpenStudy (anonymous):

3

OpenStudy (michele_laino):

finally, what is: \(8 + 3=...?\)

OpenStudy (anonymous):

11

OpenStudy (michele_laino):

correct! That is your answer!

OpenStudy (anonymous):

11

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

thx ur awesome

OpenStudy (anonymous):

thx ur awesome

OpenStudy (michele_laino):

thanks! :)